OpenStudy (anonymous):

Select the inequality the should be used to solve the following problem Jim works for a trucking company. His job is to load a truck with at least 5000 lbs of freight. He loaded 2395 lbs of freight, but has to now unload 50 lbs that was loaded by mistake. Write an inequality to solve how many more pounds he has to load. Choices below

OpenStudy (anonymous):

A)\[2395 – 50 + x \ge 5000\] B) 5000 > 2950 + x C) 5000 < 2950 + x D) x + 2395 > 5000

OpenStudy (anonymous):

2395 - 50 = 2345 so 5000 - 2345 = 2655